OpenStudy (anonymous):

The longer leg of a right triangle is 2cm less than twice the shorter leg. Find the measure of each leg of the triangle with the smallest possible area.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Jamierox4ev3r

OpenStudy (anonymous):

@gsRttN

OpenStudy (anonymous):

@Psymon

OpenStudy (anonymous):

i need help. Please?

OpenStudy (campbell_st):

why post the question again...or twice..?

OpenStudy (anonymous):

I'm sorry. No one replied with my question in this group. So i asked the question again in geometry group

OpenStudy (campbell_st):

ok... so draw a picture  find b by using pythagoras' theorem then you will find the height of the triangle and the area will be \[A = \frac{1}{2} bx\] then find the 1st derivative... \[\frac{dA}{dx}\] solve the 1st derivative for x... which will give value of x that creates the minimum area. then when you know x, you can find all 3 sides of your triangle. hope this helps

OpenStudy (anonymous):

okaay. Thank yoou so much ^_^

OpenStudy (anonymous):

the formula of finding b in Pythagorean theorem is b+ sq. root of c^2 - a^2? Am i right?

OpenStudy (anonymous):

*=

OpenStudy (campbell_st):

well b^2 = (2x -2)^2 - x^2 then take the square root. thats my best guess... I have presumed the longer side in the question is the hypotenuse. good luck... as its a difficult derivative.

OpenStudy (anonymous):

hmm. I see, well thanks for the help ^_^